OpenStudy (anonymous):

estimate the answer of 80% of 160. and explain how you got your answer.

OpenStudy (anonymous):

160 / 2 = 80. therefore 80 is 50% of 160. 80 / 2 = 40. 40+80 = 120. which is close to the exact answer of 125.

OpenStudy (anonymous):

err 128 i mean

OpenStudy (anonymous):

there might be a better way though...

OpenStudy (accessdenied):

\[ \begin{align} \frac{80}{100} &= \frac{x}{160} \qquad \text{"Cross multiply", or multiply both sides by 160*100.}\\ \\ 100x &= 12800 \\ x &= 128 \end{align} \]